OpenStudy (naga447):

I need help to Match number in standard form with its form in scientific notation. Match Term Definition 0.156 A) 1.56 × 10−1 15.6 B) 6.2 × 102 620 C) 1.56 × 101 0.062 D) 6.2 × 10−2

OpenStudy (cggurumanjunath):

1.56*10^-1=0.156

OpenStudy (cggurumanjunath):

1.56*10^+1=15.6

OpenStudy (naga447):

okay, just one question the first one that you posted "1.56*10^+1=15.6" there are not any of the options that have "10^+1" all of them are -1 or a whole number.

OpenStudy (naga447):

or 10-2

OpenStudy (cggurumanjunath):

1.56*10^+1 =1.56*10^1 =15.6 ISN'T ?

OpenStudy (cggurumanjunath):

@naga447

OpenStudy (naga447):

no the closes one to that is 1.56 × 10−1

OpenStudy (cggurumanjunath):

0.156 =1.56 × 10−1

OpenStudy (cggurumanjunath):

15.6=1.56 × 101

OpenStudy (cggurumanjunath):

620 =6.2 × 102

OpenStudy (naga447):

okay, thanks :)

OpenStudy (cggurumanjunath):

0.062= 6.2 × 10−2

OpenStudy (cggurumanjunath):

0.156 =1.56 × 10−1 BECAUSE 1.56 × 10−1 =1.56/10=0.156

OpenStudy (naga447):

Thank you!

OpenStudy (cggurumanjunath):

15.6=1.56 × 10^1 BECAUSE 1.56 × 10^1 =15.6

OpenStudy (cggurumanjunath):

:)